3 gram of activated charcoal was added to 50 ml of Acetic Acid solution in a flask after in award it was filtered and there`s and the strength of the filtrate was found to be 0.0 42 N .the amount of Acetic Acid absorb

Grade:11

1 Answers

Vikas TU
7 years ago
3 grams of charcoal that means 3/12 = 1/4 => 0.25 moles
And the molarity = 0.25*1000/50 => 5 M
Normality = Molarity/Valence factor => 5/1 = 5
Therefore, the absorbed part => 5 – 0.042 => 4.958 N
or 4.958M
Moles => 4.958*1000/50 = 99.16